A multiple-choice tesst is given with 5 choices , only one of which is correct , for each of 10 questions . What is the probability of scoring

1) exactly 60% just by guessing, 2)exactly 70% of guseeing, 3) more than 70% of guessing

how to find using excel
Answer #1

This will be a binomial distribution with parameters:

n = 10, p = P(Correct) = 1/5 = 0.20

Hence,

1) P(Exactly 60%)

= P(6 correct)

= binom.dist(6, 10, 0.20, False) [Excel Formula]

= 0.0055

2) P(Exactly 70%)

= P(7 correct)

= binom.dist(7, 10, 0.20, False) [Excel Formula]

= 0.000786

3) P(More than 70%)

= 1 - P(Less than or equal to 70%)

= 1 - binom.dist(7, 10, 0.20, True) [Excel Formula]

= 0.000078